Find the generar solution x = ( i )
Answer #1

012 my 3) () gla ( DS Then w) becomes at (pozd)-61 ;)(73) + Comb) = (*797) Day 24-12 D72 = 4+302 13 ) =0 m² 4 m+ 4 =0 arr 1